# Jacquet modules of Tate cohomology and base change lifting

### v1

@@ 4.1. Character formula for finite groups of Lie type
The part $(2)$ of the lemma is similar to part $(1)$. First assume that $(\ell,d)=1$. Note that the Frobenius $F$ on $H$ is the composition $\omega\circ{\rm Fr}$, where $\omega$ is the $d$-cycle $(1,2,\dots,d)$. Then we have the identity $F^{\ell}=\omega^{\ell}\circ{\rm Fr}^{\ell}$. Since the prime $\ell$ does not divide $d$, $w^{\ell}$ is also a $d$-cycle. Using a similar argument as in part $(1)$, we get that $S_{1}$ and $S_{2}$ are conjugate in $H^{F^{\ell}}$.
For part $(3)$ of the lemma, using the above discussion for every $(F^{\ast})^{\ell}$-conjugacy class in $W(H,S_{H})$, we can associate $\ell$-partitions $(r_{i1},r_{i2},\dots,r_{ik_{i}})$ of the positive integer $d/\ell$; and the torus associated with this $(F^{\ast})^{\ell}$-conjugacy class of $W(H,S_{H})$ is ${\rm GL}_{n}(\mathbb{F}_{q^{\ell}})$-conjugate to a torus of ${\rm GL}_{n}$ corresponding to the partition
@@ 4.2. Shintani correspondence
We denote by $\sigma$ the Frobenius element in ${\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q})$, i.e., $\sigma(x)=x^{q}$ for $x\in\mathbb{F}_{q^{n}}$. Let $\Pi$ be an irreducible $\ell$-adic representation of ${\rm GL}_{n}(\mathbb{F}_{q^{\ell}})$ such that $\Pi$ is ${\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q})$ equivariant, i.e., $\Pi\simeq\Pi^{\sigma}$. In the article [Shi76, Theorem 1], Shintani proved that for a suitable choice of isomorphism $I_{\sigma}:\Pi\rightarrow\Pi^{\sigma}$, there exists an irreducible representation $\pi$ of ${\rm GL}_{n}(\mathbb{F}_{q})$ such that
${\rm Tr}(I_{\sigma}\circ\Pi(g))={\rm Tr}\big{(}\pi({\rm Norm}_{\mathbb{F}_{q^{l}}/\mathbb{F}_{q}}(g))\big{)},\ \forall g\in{\rm GL}_{n}(\mathbb{F}_{q^{l}}),$ (4.3)
where ${\rm Norm}_{\mathbb{F}_{q^{l}}/\mathbb{F}_{q}}(g)$ is the unique conjugacy class of ${\rm GL}_{n}(\mathbb{F}_{q})$ containing the element
@@
The association $\Pi\mapsto\pi$ gives a bijection from the set of all isomorphism classes of $\sigma$ equivariant irreducible $\ell$-adic representation of ${\rm GL}_{n}(\mathbb{F}_{q^{\ell}})$ onto the set of isomorphism classes of irreducible $l$-adic representations of ${\rm GL}_{n}(\mathbb{F}_{q})$. The inverse of this bijection is called the Shintani base change lifting, denoted by ${\rm bc}_{\mathbb{F}_{q^{l}}/\mathbb{F}_{q}}$.
@@ 4.3. Congruence of characters
In this subsection, we prove certain congruences between the character of a cuspidal representation and the character of its Shintani lifting to ${\rm GL}_{n}(\mathbb{F}_{q^{\ell}})$, when evaluated at an element of ${\rm GL}_{n}(\mathbb{F}_{q})$. We follow the notations of the preceding subsection 4.2, and let $F$ be the Frobenius endomorphism of ${\rm GL}_{n}/\mathbb{F}_{q}$ induced by the Frobenius automorphism $\overline{\mathbb{F}}_{q}\rightarrow\overline{\mathbb{F}}_{q}$,  $x\mapsto x^{q}$. We begin with the following preliminary lemma. For any representation $\tau$ of ${\rm GL}_{n}(\mathbb{F}_{q^{\ell}})$, we denote by $\tau^{F}$ the composition of the representation $\tau$ with the automorphism of ${\rm GL}_{n}(\mathbb{F}_{q^{\ell}})$ induced by $\sigma$.
Lemma 4.2.
@@
Recall that $[w_{0}^{r}({\rm id}+w)](i)=(i+r)+w(i+r)$, $10\leq r\leq\ell-$ 1, and we have
@@
Proposition 4.3.
Let $n=\ell m$. Let $\Pi$ be an absolutely irreducible generic $\mathcal{K}$ representation of ${\rm GL}_{n}(\mathbb{F}_{q^{\ell}})$ such that $\Pi\simeq\Pi^{\sigma}$, for all $\sigma\in{\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q})$. Let $\pi$ be the $\ell$-adic cuspidal representation of ${\rm GL}_{n}(\mathbb{F}_{q})$ such that $\Pi\otimes_{\mathcal{K}}\overline{\mathbb{Q}}_{\ell}$ is the Shintani lift of $\pi$. Let $\mathcal{L}$ be a $\Lambda[{\rm GL}_{n}(\mathbb{F}_{q^{\ell}})\rtimes{\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q})]$ stable lattice in $\Pi$. Then we have
@@
Let $\sigma$ be the Frobenius element in ${\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q})$, and the Frobenius endomorphism of ${\rm GL}_{n}/\mathbb{F}_{q}$ is denoted by $F$. We fix a ${\rm GL}_{n}(\mathbb{F}_{q^{l}})\rtimes\langle\sigma\rangle$ stable $\Lambda$-lattice $\mathcal{L}$ in $\Pi$. Since the Tate cohomology group $\widehat{H}^{0}({\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q}),\mathcal{L})$ is irreducible and cuspidal (Theorem 3.6), it is sufficient to show that
for all $g$ with Jordan decomposition of the form $su$, where the characteristic polynomial of $s$ is of the form $f(X)^{r}$ for some irreducible polynomial $f(X)\in\mathbb{F}_{q}[X]$; let $\mathcal{S}$ be the set of such elements $g\in{\rm GL}_{n}(\mathbb{F}_{q})$. Since the Tate cohomology group $\widehat{H}^{1}({\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q}),\mathcal{L})$ is trivial (Theorem 3.6), it follows from Lemma 2.4 that the above assertion is equivalent to the following identity
@@
Let $S_{0}$ be an $F$-stable split maximal torus of ${\rm GL}_{n}$. Let $T$ be an $F$-stable maximal torus of ${\rm GL}_{n}$ such that $(T,F)$ is isomorphic to $(S_{0},{\rm Ad}(w_{0})\circ F)$, where $w_{0}$ is an $n$-cycle in $W({\rm GL}_{n},S_{0})$. Let $\theta:T^{F}\rightarrow\mathcal{K}^{\ast}$ be a character in general position such that $\pi=R_{T,\theta}$. Let ${\rm Nr}:T^{F^{\ell}}\rightarrow T^{F}$ be the norm map
@@
The Shintani lifting of $\pi$ to ${\rm GL}_{n}(\mathbb{F}_{q^{\ell}})$ is equal to $\Pi=R^{(\ell)}_{T,\eta}$. Here $R^{(\ell)}_{T,\eta}$ is the Deligne–Lusztig induction of the character $\eta$ of $T^{F^{\ell}}$ to $G^{F^{\ell}}$. Note that $T^{F^{\ell}}$ is identified with $(\mathbb{F}_{q^{n}}^{\ast})^{\ell}$ and $T^{F}$ is identified with $\mathbb{F}_{q^{n}}^{\ast}$ such that ${\rm Nr}$ is given by
@@
Let $P$ be an $F^{\ell}$-stable parabolic subgroup containing $T$, and let $M$ be an $F^{\ell}$-stable Levi subgroup of $P$ containing $T$. We can identify the group $M^{F^{\ell}}$ with ${\rm GL}_{n/\ell}(\mathbb{F}_{q^{\ell}})^{\ell}$ and $R^{(\ell)}_{T,\eta}$ with the parabolic induction
@@
where $\tau$ is the cuspidal representation of ${\rm GL}_{n/\ell}(\mathbb{F}_{q^{\ell}})$ obtained as the Deligne–Lusztig induction of the character $\theta$ of $\mathbb{F}_{q^{n}}^{\ast}$ (which is considered as $F^{\ell}$ fixed points of the $F^{\ell}$-stable maximal torus of ${\rm GL}_{n/\ell}/\mathbb{F}_{q^{\ell}}$ corresponding to the $n/\ell$-cycle).
Let $s\in T^{F}$ be an element with characteristic polynomial $f(X)^{r}$, where $f(X)\in\mathbb{F}_{q}[X]$ is an irreducible polynomial of degree $d$. We have $dr=n$. Let $H$ be the group $C_{{\rm GL}_{n}}(s)$. Let $u\in H^{F}$ be an unipotent element. From the Deligne–Lusztig character formula (4.1), we get that
$\chi_{\Pi}(su)=Q_{H,T,F^{\ell}}(u)\sum_{w\in W(G,T)^{F^{\ell}}/W(H,T)^{F^{\ell}}}\eta(wsw^{-1})$
@@
Note that $H^{F^{\ell}}$ con be identified with ${\rm GL}_{r}(\mathbb{F}_{q^{d}})^{\ell}$ and the element $u\in H^{F}\subset H^{F^{\ell}}$ with
@@
Using Lemma 4.2, we get that
@@
Using Lemma 4.2, we get that
@@
Assume that $\mu=(\mu_{1},\mu_{2},\dots,\mu_{k})$ is the partition associated to the conjugacy class of $u$ in $H^{F}$. Then $\mu$ is the partition of the $H^{F^{\ell}}$ conjugacy class of $u$. The partition associated with $T^{F}$ is $(n)$ and $(n/\ell,n/\ell,\dots,n/\ell)$ is the partition of $T^{F^{\ell}}$. Let $w\in S_{n}$ be an $n$-cycle. It follows from Springer that
@@
This concludes the proof of the proposition. ∎
@@
Theorem 4.4.
Let $\Pi$ be an absolutely irreducible generic $\mathcal{K}$-representation of ${\rm GL}_{n}(\mathbb{F}_{q^{\ell}})$. Let $\pi$ be an irreducible $\ell$-adic cuspidal representation of ${\rm GL}_{n}(\mathbb{F}_{q})$ such that $\Pi\otimes_{\mathcal{K}}\overline{\mathbb{Q}}_{\ell}$ is the Shintani lift of $\pi$. Then for any ${\rm GL}_{n}(\mathbb{F}_{q^{\ell}})\rtimes{\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q})$ stable $\Lambda$-lattice $\mathcal{L}$ in $\Pi$, the Tate cohomology group $\widehat{H}^{0}({\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q}),\mathcal{L})$ is isomorphic to the Frobenius twist $r_{\ell}(\pi)^{(\ell)}$.
@@
Let $\sigma$ be the Frobenius element of ${\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q})$. Note that the Shintani lift $\Pi$ is either cuspidal or a parabolically induced representation of the form
$\tau\times\tau^{\sigma}\times\cdots\times\tau^{\sigma^{\ell-1}}$ (4.7)
where $\tau$ is an irreducible cuspidal $\mathcal{K}$-representation of ${\rm GL}_{m}(\mathbb{F}_{q^{\ell}})$, $n=m\ell$, such that $\tau$ is not isomorphic to $\tau^{\sigma}$. When $\Pi$ is a cuspidal representation, then the theorem follows from [Ron16, Theorem 13]. If $\Pi$ is of the form (4.7), the theorem follows from Proposition 4.3. ∎

### v2

@@ 5.1. Character formula for finite groups of Lie type
The part (2)(2)( 2 ) of the lemma is similar to part (1)(1)( 1 ). First assume that $(\ell,d)=1$. Note that the Frobenius endomorphism $F$ on $H$ is the composition $\omega\circ{\rm Fr}$, where $\omega$ is the $d$-cycle $(1,2,\dots,d)$. Then we have the identity $F^{\ell}=\omega^{\ell}\circ{\rm Fr}^{\ell}$. Since $\ell$ does not divide $d$, $w^{\ell}$ is also a $d$-cycle. Using a similar argument as in part (1)(1)( 1 ), we get that $S_{1}$ and $S_{2}$ are conjugate in $H^{F^{\ell}}$.
For part (3)(3)( 3 ) of the lemma, using the above discussion for every $(F^{\ast})^{\ell}$-conjugacy class in $W(H,S_{H})$, we can associate $\ell$-partitions $(r_{i1},r_{i2},\dots,r_{ik_{i}})$ of the positive integer $d/\ell$, for $1\leq i\leq\ell$; and the torus associated with this $(F^{\ast})^{\ell}$-conjugacy class of $W(H,S_{H})$ is ${\rm GL}_{n}(\mathbb{F}_{q^{\ell}})$-conjugate to a torus of ${\rm GL}_{n}$ corresponding to the partition
@@ 5.2. Shintani correspondence
Let $\sigma$ be the Frobenius element in ${\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q})$. Let $\Pi$ be an $\ell$-adic irreducible representation of ${\rm GL}_{n}(\mathbb{F}_{q^{\ell}})$ such that $\Pi$ is ${\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q})$ equivariant, i.e., $\Pi\simeq\Pi^{\sigma}$. In the article [Shi76, Theorem 1], Shintani proved that for a suitable choice of isomorphism $I_{\sigma}:\Pi\rightarrow\Pi^{\sigma}$, there exists an $\ell$-adic irreducible representation $\pi$ of ${\rm GL}_{n}(\mathbb{F}_{q})$ such that
${\rm Tr}(I_{\sigma}\circ\Pi(g))={\rm Tr}\big{(}\pi({\rm Norm}_{\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q}}(g))\big{)},\ \forall g\in{\rm GL}_{n}(\mathbb{F}_{q^{\ell}}),$ (5.3)
where ${\rm Norm}_{\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q}}(g)$ is the unique conjugacy class of ${\rm GL}_{n}(\mathbb{F}_{q})$ containing the element
@@
The association $\Pi\mapsto\pi$ gives a bijection from the set of all isomorphism classes of ${\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q})$ equivariant irreducible $\ell$-adic representation of ${\rm GL}_{n}(\mathbb{F}_{q^{\ell}})$ onto the set of isomorphism classes of irreducible $\ell$-adic representations of ${\rm GL}_{n}(\mathbb{F}_{q})$. The inverse of this bijection is called the Shintani base change lifting, denoted by ${\rm bc}_{\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q}}$.
@@ 5.3. Congruence of characters
In this subsection, we prove certain congruences between the character of a cuspidal representation and the character of its Shintani lifting to ${\rm GL}_{n}(\mathbb{F}_{q^{\ell}})$, when evaluated at an element of ${\rm GL}_{n}(\mathbb{F}_{q})$. We follow the notations of the preceding subsection 5.2, and let $F$ be the Frobenius endomorphism of ${\rm GL}_{n}/\mathbb{F}_{q}$ induced by the Frobenius automorphism $\overline{\mathbb{F}}_{q}\rightarrow\overline{\mathbb{F}}_{q}$,  $x\mapsto x^{q}$. We begin with the following preliminary lemma. For any representation $\tau$ of ${\rm GL}_{n}(\mathbb{F}_{q^{\ell}})$, we denote by $\tau^{F}$ the composition of the representation $\tau$ with the automorphism of ${\rm GL}_{n}(\mathbb{F}_{q^{\ell}})$ induced by $\sigma$. For any representation $\rho$, we denote by $\chi_{\rho}$ its character.
Lemma 5.3.1.
@@
Recall that $[w_{0}^{r}({\rm id}+w)](i)=(i+r)+w(i+r)$, $0\leq r\leq\ell-1$, and we have
@@
5.4.
Let $G\rtimes\Gamma$ be a finite group such that $\Gamma\simeq\mathbb{Z}/\ell\mathbb{Z}$, and we fix a generator $\sigma$ of $\Gamma$. The following lemma, in certain cases, gives the trace of $G^{\sigma}$ action on the Tate cohomology with respect to the action of $\Gamma$. This will be useful when we consider the Tate cohomology of a lattice in Shintani lifts.
Lemma 5.4.1.
Let $G$ be a finite group with an automorphism $\sigma$ of order $\ell$. Let $\Gamma$ be the group generated by $\sigma$. Let $M$ be a $\Lambda[G\rtimes\Gamma]$-module such that $M$ is a free $\Lambda$-module of finite rank. Assume that $\widehat{H}^{1}(\Gamma,M)=0$. Then
${\rm Tr}(g|\widehat{H}^{0}(\Gamma,M))=\overline{{\rm Tr}(g|M)},$
for all $g\in G^{\sigma}$.
Proof.
This lemma was observed by Ronchetti ([Ron16, Theorem 12]). Since, this lemma is used in a crucial way, we give a proof. We may assume that there is a finite unramified extension $K$ of $\mathbb{Q}_{\ell}$ and a $\mathfrak{o}_{K}[G\rtimes\Gamma]$ module $M_{0}$ such that $M_{0}$ is a free $\mathfrak{o}_{K}$ module and $M=M_{0}\otimes_{\mathfrak{o}_{K}}\Lambda$. Assume that $M_{0}=\oplus_{i=1}^{k}N_{i}$, where $N_{i}$ is an indecomposable $\mathbb{Z}_{\ell}[\Gamma]$ submodule of $M$. Since $\widehat{H}^{1}(\Gamma,M)=0$, we get that $N_{i}$ is isomorphic to $\mathbb{Z}_{\ell}$ or $\mathbb{Z}_{\ell}[\Gamma]$. Let $p_{i}:N_{i}\rightarrow M$ and $q_{i}:M\rightarrow N_{i}$ be the embedding and the projection maps, respectively. Note that
${\rm Tr}(g|M)=\sum_{i=1}^{k}{\rm Tr}(q_{i}gp_{i}:N_{i}\rightarrow N_{i}).$
The action of $q_{i}gp_{i}$ on $N_{i}$ has a unique eigenvalue. When $N_{i}=\mathbb{Z}_{\ell}$, the element $q_{i}gp_{i}$ acts by an unit $\xi\in\mathbb{Z}_{\ell}^{\ast}$ whose $-\ell$ reduction $\overline{\xi}$ coincides with the action of $\xi$ on the Tate cohomology space $\widehat{H}^{0}(\Gamma,\mathbb{Z}_{\ell})=\mathbb{Z}/\ell\mathbb{Z}$. Let us consider the case where $N_{i}=\mathbb{Z}_{\ell}[\Gamma]$. Note that $\widehat{H}^{0}(\Gamma,N_{i})=0$. If $\lambda$ is the unique eigenvalue of $q_{i}gp_{i}$ on ${N_{i}}$, then ${\rm Tr}(q_{i}gp_{i})=\ell.\lambda$. Thus,
$\overline{{\rm Tr}(g|_{M})}=0={\rm Tr}(g|_{\widehat{H}^{0}(\Gamma,M)}).$
Hence the lemma. ∎
Proposition 5.4.2.
Let $\Pi$ be an absolutely irreducible generic $\mathcal{K}$ representation of ${\rm GL}_{n}(\mathbb{F}_{q^{\ell}})$ such that $\Pi\simeq\Pi^{\sigma}$, for all $\sigma\in{\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q})$. Let $\pi$ be the $\ell$-adic cuspidal representation of ${\rm GL}_{n}(\mathbb{F}_{q})$ such that $\Pi\otimes_{\mathcal{K}}\overline{\mathbb{Q}}_{\ell}$ is the Shintani lift of $\pi$. Let $\mathcal{L}$ be a $\Lambda[{\rm GL}_{n}(\mathbb{F}_{q^{\ell}})\rtimes{\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q})]$ stable lattice in $\Pi$. Then we have
@@
Let $\sigma$ be the Frobenius element in ${\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q})$ and let $F$ be the corresponding Frobenius endomorphism of ${\rm GL}_{n}$. We fix a ${\rm GL}_{n}(\mathbb{F}_{q^{\ell}})\rtimes\langle F\rangle$ stable $\Lambda$-lattice $\mathcal{L}$ in $\Pi$. Since the Tate cohomology group $\widehat{H}^{0}({\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q}),\mathcal{L})$ is irreducible and cuspidal (Theorem 4.1.1), it is sufficient to show that
${\rm Tr}(g|{\widehat{H}^{0}({\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q}),\mathcal{L})})=\overline{\chi_{\pi}(g)}^{\ell},$ (5.7)
for all $g$ with Jordan decomposition $g=su$, where the characteristic polynomial of $s$ is of the form $f(X)^{r}$ for some irreducible polynomial $f(X)\in\mathbb{F}_{q}[X]$; let $\mathcal{S}$ be the set of all such elements $g\in{\rm GL}_{n}(\mathbb{F}_{q})$. Since the Tate cohomology group $\widehat{H}^{1}({\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q}),\mathcal{L})$ is trivial (Theorem 4.1.1), it follows from Lemma 5.4.1 that the equality (5.7) is equivalent to the following identity
@@
[…]
@@
The Shintani lifting of $\pi$ to ${\rm GL}_{n}(\mathbb{F}_{q^{\ell}})$ is equal to $\Pi=(-1)^{n-\ell}R^{(\ell)}_{T,\eta}$ (see [Kaw87] for a reference). Here $R^{(\ell)}_{T,\eta}$ is the Deligne–Lusztig induction of the character $\eta$ of $T^{F^{\ell}}$ to $G^{F^{\ell}}$.
First we assume that $n=\ell m$ for some integer $m$. Note that $T^{F^{\ell}}$ is identified with $(\mathbb{F}_{q^{n}}^{\ast})^{\ell}$ and $T^{F}$ is identified with $\mathbb{F}_{q^{n}}^{\ast}$ such that ${\rm Nr}$ is given by
@@
Let $P$ be an $F^{\ell}$-stable parabolic subgroup containing $T$, and let $M$ be an $F^{\ell}$-stable Levi subgroup of $P$ containing $T$. We can identify the group $M^{F^{\ell}}$ with ${\rm GL}_{n/\ell}(\mathbb{F}_{q^{\ell}})^{\ell}$ and $\Pi=(-1)^{n-\ell}R^{(\ell)}_{T,\eta}$ with the parabolic induction
@@
where $\tau$ is the irreducible cuspidal representation of ${\rm GL}_{n/\ell}(\mathbb{F}_{q^{\ell}})$, obtained as the Deligne–Lusztig induction of the character $\theta$ of $\mathbb{F}_{q^{n}}^{\ast}$ (which is considered as $F^{\ell}$-fixed points of a $F^{\ell}$-stable maximal torus of ${\rm GL}_{n/\ell}$ corresponding to an $n/\ell$-cycle).
Let $s\in T^{F}$ be an element with characteristic polynomial $f(X)^{r}$, where $f(X)\in\mathbb{F}_{q}[X]$ is an irreducible polynomial of degree $d$. We have $n=dr$. Let $H$ be the centralizer $C_{{\rm GL}_{n}}(s)$. Let $u\in H^{F}$ be an unipotent element. From the Deligne–Lusztig character formula (5.1) and Lemma 5.1.1, we get that
@@
Note that $H^{F^{\ell}}$ can be identified with ${\rm GL}_{r}(\mathbb{F}_{q^{d}})^{\ell}$ and the element $u\in H^{F}\subset H^{F^{\ell}}$ with
@@
Using Lemma 5.3.1, we get that
@@
Using Lemma 5.3.1, we get that
@@
[…]
@@
In the case where $(\ell,n)=1$, the above proposition is proved by Ronchetti ([Ron16, Theorem 13]). For the uniformity of the proof, we treat this case using our method. For $g\in\mathcal{S}$, we have
$R_{T,\eta}^{(\ell)}(g)=Q_{H,T,F^{\ell}}(u)\sum_{w\in W(G,T)^{F^{\ell}}/W(H,T)^{F^{\ell}}}\theta(wsw^{-1})$
and in this case, we have
$H^{F}={\rm GL}_{r}(\mathbb{F}_{q}),\ H^{F^{\ell}}={\rm GL}_{r}(\mathbb{F}_{q^{d\ell}}),\ T^{F}=\mathbb{F}_{q^{\ell}},\ T^{F^{\ell}}=\mathbb{F}_{q^{n\ell}}$
and we have
$Q_{H,T,F^{\ell}}(u)=\sum_{i\geq 0}a_{i}(w,u)q^{i\ell}\equiv(\sum_{i\geq 0}a_{i}(w,u)q^{i})^{\ell}\ \text{mod}\ (\ell).$
Then we have
$\sum_{w\in W(G,T)^{F^{\ell}}/W(H,T)^{F^{\ell}}}\theta(wsw^{-1})=\sum_{i=1}^{d}\theta(s^{q^{i\ell}})\equiv[\sum_{i=1}^{d}\theta(s^{q^{i}})]^{\ell}\ \text{mod}\ (\ell)$
@@
Theorem 5.4.3.
Let $\Pi$ be an absolutely irreducible generic $\mathcal{K}$-representation of ${\rm GL}_{n}(\mathbb{F}_{q^{\ell}})$. Let $\pi$ be an irreducible $\ell$-adic cuspidal representation of ${\rm GL}_{n}(\mathbb{F}_{q})$ such that $\Pi\otimes_{\mathcal{K}}\overline{\mathbb{Q}}_{\ell}$ is the Shintani lift of $\pi$. Then for any ${\rm GL}_{n}(\mathbb{F}_{q^{\ell}})\rtimes{\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q})$ stable $\Lambda$-lattice $\mathcal{L}$ in $\Pi$, the Tate cohomology $\widehat{H}^{0}({\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q}),\mathcal{L})$ is isomorphic to the Frobenius twist $r_{\ell}(\pi)^{(\ell)}$.
@@
To prove the theorem, we use the fact that two irreducible $\ell$-modular representations of a finite group are isomorphic if and only if they have the same character ([Lam01, Corollary 7.21]). Using Proposition 5.4.2, we get the character identity
${\rm Tr}(g|_{\widehat{H}^{0}({\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q}),\mathcal{L})})=\overline{\chi_{\pi}(g)}^{\ell},$
for all $g\in{\rm GL}_{n}(\mathbb{F}_{q})$. Note that the Tate cohomology group $\widehat{H}^{0}({\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q}),\mathcal{L})$ and the Frobenius twist $r_{\ell}(\pi)^{(\ell)}$ are both irreducible. Thus, the Tate cohomology $\widehat{H}^{0}({\rm Gal}(\mathbb{F}_{q^{\ell}}/\mathbb{F}_{q}),\mathcal{L})$ is isomorphic to $r_{\ell}(\pi)^{(\ell)}$ as ${\rm GL}_{n}(\mathbb{F}_{q})$ representations. Hence the theorem. ∎